Suppose that 48% of the women who gave birth at a certain hospital last year were over 30 years old, and that 46% were unmarried. If 65% of the women were over 30 or unmarried (or both), what is the probability that a woman who gave birth at the hospital was both unmarried and over 30 ?

Answer by stanbon(48569)

Suppose that 48% of the women who gave birth at a certain hospital last year were over 30 years old, and that 46% were unmarried. If 65% of the women were over 30 or unmarried (or both), what is the probability that a woman who gave birth at the hospital was both unmarried and over 30 ?
----------------------------
Formula: P(A and B) = P(A) + P(B) - P(A or B)
-----------------------------
Your problem:
P(un and >30) = P(un) + P(>30) - P(un or >30)
P(un and >30) = 0.46 + 0.48 - 0.65 = 0.29
==========================
Cheers,
Stan H.